OpenStudy (anonymous):

integrate sinhx/cosh^2 x dx, do i have to do substitution here

zepdrix (zepdrix):

Yes.

zepdrix (zepdrix):

The most natural sub is either going to be u=sinh(x) or u=cosh(x). Any ideas which would be better? :)

zepdrix (zepdrix):

I guess I should say, Any ideas which sub `will actually work`? :)

OpenStudy (anonymous):

u=coshx?

zepdrix (zepdrix):

\[\large\rm \int\limits\frac{\sinh x}{\cosh^2x}dx\qquad=\qquad \int\limits \frac{1}{\color{royalblue}{\cosh}^2\color{royalblue}{x}}\left(\color{orangered}{\sinh x~dx}\right)\]Good :)

zepdrix (zepdrix):

\[\large\rm \color{royalblue}{u=\cosh x},\qquad\qquad du=?\]

OpenStudy (anonymous):

sinhx

zepdrix (zepdrix):

\[\large\rm \color{royalblue}{u=\cosh x},\qquad\qquad \color{orangered}{du=\sinh x~dx}\]Cool

OpenStudy (anonymous):

so it will leave \[\int\limits \frac{ 1 }{ coshx }\]

zepdrix (zepdrix):

Woops, replace all of the colors,\[\large\rm \int\limits\limits \frac{1}{\color{royalblue}{\cosh}^2\color{royalblue}{x}}\left(\color{orangered}{\sinh x~dx}\right)\quad=\quad \int\limits\limits \frac{1}{\color{royalblue}{u}^2}\left(\color{orangered}{du}\right)\]Understand?

OpenStudy (anonymous):

yup

zepdrix (zepdrix):

And then just power rule through your u, ya?\[\large\rm =\int\limits u^{-2}du\]

OpenStudy (anonymous):

ok, thanks